OpenStudy (anonymous):

write in slope-intercept form. passes through (2,4) and (5,4).

OpenStudy (lasttccasey):

There is a very useful formula for this, two actually. (y-yi) = m(x-xi) where m = (y1-y2)/(x1-x2) or rise/run

OpenStudy (anonymous):

I came up with m=0. can I still write a slope intercept equation

OpenStudy (lasttccasey):

Well if the slope was zero, it would be a straight line essentially. In this case it will be zero so you will get something like y = #